# Two point functions and quantum fields in the anti-de Sitter universe [^1] Dedicated to the memory of Henri Epstein

### v1

@@ 1 Introduction
$W^{(d)}_{\lambda}(z_{1},z_{2})=c_{d}(\lambda)\int_{\gamma(z_{1})}(z_{1}\cdot\zeta)^{\lambda}(z_{2}\cdot\zeta)^{-\lambda-d+1}d\mu_{\gamma}(\zeta).$ (1.1)
@@
The significance of the new representation (1.1) is evident: it provides the closest possible analogue, within the AdS framework, to the plane-wave Fourier representation of the two-point function for a massive scalar field in (complex) Minkowski space $M_{d}^{(c)}$ with positive energy spectrum [12]:
@@
Manifestly covariant momentum space representations of quantum fields provide the physically intuitive ingredients of Feynman diagrammatics [13] and a great deal of the predictive power of quantum field theory depends on them. Furthermore, many crucial mathematical developments such as the PCT and the Spin-Statistics theorems [12] and the Euclidean formulation of quantum field theory [14] are based on the plane-wave Fourier representation of the correlation functions and the positivity of the energy spectrum. The importance of the Fourier representation (1.2) and of the corresponding representation of the Feynman propagator cannot be overestimated and we expect that (1.1) to occupy the same position in AdS QFT.
The second major topic addressed in this work pertains to the horospherical, or Poincaré, coordinate system, which covers exactly one half of the AdS manifold. We analyze the Fourier transform of the plane waves w.r.t. the Poincaré coordinates and show how this transform diagonalizes the two-point function (1.1). We present the final, remarkably elegant result here deferring detailed explanations to Sections 13 and 14:
@@
$=\frac{\left(uu^{\prime}\right)^{\frac{d-1}{2}}}{2(2\pi)^{d-2}}\int e^{-i{\rm k}({\rm z}_{1}-{\rm z}_{2})}\theta({\rm k}^{0})\theta({\rm k}^{2})J_{\frac{d-1}{2}+\lambda}\left(u\sqrt{{\rm k}^{2}}\right)J_{\frac{d-1}{2}+\lambda}\left(u^{\prime}\sqrt{{\rm k}^{2}}\right)d{\rm k}$ (1.4)
The crucial role played by the global analyticity properties of the plane waves, which are inherited from our foundational axioms, will appear clearly in this calculation.
Moreover, this analysis yields an unexpected insight: Euclidean Feynman diagrams constructed on Lobachevski space can be mapped, via Wick rotation, onto diagrams entirely supported in the Poincaré patch. Multiple examples are given that substantiate the conjecture that any Feynman or Witten diagram computed by integration solely over a Poincaré patch of the real AdS manifold remains AdS covariant, even though the patch itself is not invariant.
@@ 2 AdS and its covering: geometrical setup
$G_{0}$ acts transitively on $AdS_{d}$; we chose
@@
to be the distinguished origin of $AdS_{d}$; in the folowing we will take for simplicity a unit radius $R=1$. The restriction to $AdS_{d}$ of the pseudo-Riemannian metric $g_{\mu\nu}dx^{\mu}dx^{\nu}$ is a Lorentzian metric with signature (+,,,)(+,\ -$\ldots,$-). The complex $d$-dimensional anti-de Sitter manifold $AdS_{d}^{(c)}$ can analogously be defined by its embedding in the complex space ${\bf C}^{d+1}_{2}$ endowed with the metric $g_{\mu\nu}$:
$AdS_{d}^{(c)}=\{z=x+iy\in{\bf C}^{d+1}_{2},\ z^{2}=1\}=\{z=x+iy,\ x^{2}-y^{2}=1,\ x\cdot y=0\}.$ (2.9)
@@
It is possible to single out in $AdS_{d}^{(c)}$ a Euclidean manifold $EAdS_{d}$; we choose the connected real submanifold defined by taking $z^{0}=ix^{0},\ x^{1},\ldots,x^{d}$ real, $x^{d}>0$:
$z^{\scriptscriptstyle{E}}=(ix^{0},\vec{x},x^{d}).$ (2.10)
It is the upper sheet of the two-sheeted hyperboloid with equation
$(x^{d})^{2}-(x^{0})^{2}-{\vec{x}}^{2}=1,\ \ \ x^{d}\geq 1$ (2.11)
equipped with the Riemannian metric induced by the scalar product (2.5). Global coordinates may be introduced on $AdS_{d}$ by the map
$(t,\,\vec{x})\mapsto\left(\sqrt{1+\vec{x}^{2}}\,\sin t,\ \vec{x},\ \sqrt{1+\vec{x}^{2}}\,\cos t\right)$ (2.12)
@@
$(s,\,\vec{x})\mapsto(i\sqrt{1+\vec{x}^{2}}\,{\rm sh}\ s,\ \vec{x},\ \sqrt{1+\vec{x}^{2}}\,{\rm ch}\ s).$ (2.13)
@@ 4 The chiral tuboids $\cal Z$$_{+}$ and $\cal Z$$_{-}$
(ii) The image of either of the domains $\cal Z$$_{+}$ or $\cal Z$$_{-}$ under the coordinate map $z\mapsto z^{d}=e_{d}\cdot z$ is the cut-plane $\Delta={\bf C}\setminus[-1,1]$.
(iii) The image of the domain ${\cal Z}_{-}\times{\cal Z}_{+}$ (or ${\cal Z}_{+}\times{\cal Z}_{-}$) by the scalar product mapping $z_{1},z_{2}\mapsto z_{1}\cdot z_{2}$ is the cut-plane $\Delta$.
(iv) The map $z_{1},z_{2}\mapsto z_{1}\cdot z_{2}$ of ${\cal Z}_{-}\times{\cal Z}_{+}$ onto $\Delta$ can be lifted to a map of $\widetilde{\cal Z}_{-}\times\widetilde{\cal Z}_{+}$ onto the covering $\widetilde{\Delta}$ of the cut-plane $\Delta$.
@@ 2. Locality: if $x_{1}$ and $x_{2}$ are space-like separated then the commutator
$4^{\prime\prime}$. Maximal analyticity property: the reduced two-point function $W(z_{1}\cdot z_{2})$ extends to a function analytic in the covering $\widetilde{\Delta}$ of the cut-plane $\Delta=\{{\bf C}\setminus[-1,1]\}.$ For theories single-valued on $AdS_{d}$, $W(z_{1}\cdot z_{2})$ is analytic in $\Delta$.
@@ 6 Klein-Gordon equation, homogeneous functions and plane waves
This remains true if $h$ is a $G_{0}$-invariant distribution such as $\delta((x\cdot x)-1)$. In particular the operators ${\cal M}_{\mu\nu}$ and ${\cal M}^{2}$ commute with the restriction to invariant submanifolds of ${\bf R}^{d+1}_{2}$ and also with dilations. If $f$ is a smooth function or a distribution homogeneous of degree 0 on the open set $\Omega=\{x\in{\bf R}^{d+1}_{2}\ :\ (x,x)>0\}$ then
@@
Eqs. (6.47) and (6.48) then imply that the restriction of $f$ to the anti de Sitter manifold $AdS_{d}$ satisfies the Klein-Gordon equation
$(\Box+\lambda(\lambda+d-1))(f(x)|_{AdS_{d}})=0$ (6.49)
@@
For real $\lambda$ and real $d$ the squared mass attains a minimum value which is negative:
$\lambda_{d}=-\frac{d-1}{2},\ \ \ \ m^{2}_{\lambda_{d}}=-\frac{(d-1)^{2}}{4}.$ (6.51)
It may be useful to introduce the complex parameter $\nu=\nu_{\lambda}$ by subtracting $\lambda_{d}$:
$\nu=\lambda+\frac{d-1}{2}$ (6.52)
so that $\nu=0$ corresponds precisely to the minimal value of the squared mass
$m_{\lambda}^{2}=\nu_{\lambda}^{2}-\frac{(d-1)^{2}}{4}.$ (6.53)
In the following we will use $\lambda$ and $\nu$ interchangeably.
Plane waves are the most important example of homogeneous solution of the massless equation in the ambient space:
$x,\zeta\mapsto(x\cdot\zeta)^{\lambda}$ (6.54)
where $\lambda$ is a complex number and $\zeta$ is a complex null vector of ${\cal C}_{d}^{(c)}$. Taking the restrictions to the anti de Sitter universe
$\phi_{\lambda}(x,\zeta)=(x\cdot\zeta)^{\lambda},\ \ \ x\in AdS_{d}$ (6.55)
satisfies the massive Klein-Gordon equation which we rewrite for clarity:
$(\Box+m_{\lambda}^{2})\phi_{\lambda}(x,\zeta)=0.$ (6.56)
The invariance of $m^{2}_{\lambda}$ under the involution
$\lambda\to 1-d-\lambda$ (6.57)
implies that, given the complex null vector $\zeta$, there is a second solution of Eq. (6.56) obtained by applying the involution to the exponent of the wave:
@@
Now the main difficulty arises: although the complex parameter $\lambda$ in the above definition is unrestricted, the plane waves themselves are only locally well-defined on $AdS_{d}$. An exception occurs when $\lambda$ takes integer values; in that case the plane waves are globally well-defined. In the general case, however, it requires additional effort to construct single-valued plane waves on the covering space $\widetilde{}AdS_{d}$. This issue will be addressed in detail in the next section.
@@ 7 Chiral cones
    The image of either $\cal C$$_{+}$ or $\cal C$$_{-}$ under the coordinate map $\zeta\mapsto\zeta^{d}=e_{d}\cdot\zeta$ is the punctured complex-plane ${\bf C}^{*}={\bf C}\setminus\{0\}$. This fact is intuitively clear as the chiral cones are the limit of the chiral tuboids when the radius $R$ tends to zero.
@@
    Eqs. (7.64) and (7.66) then imply that the image of ${\cal Z}_{-}\times\overline{{\cal C}}_{+}$ by the scalar product mapping $z,\zeta\mapsto z\cdot\zeta$ is also the punctured complex-plane $\bf C$$^{*}$ and the same is obviously true for ${\cal Z}_{+}\times\overline{{\cal C}}_{-}$. In words: the scalar product $z\cdot\zeta$ cannot vanish when either $z\in{\cal Z}_{-}$ and $\zeta\in\overline{{\cal C}}_{+}$ or when $z\in{\cal Z}_{+}$ and $\zeta\in\overline{{\cal C}}_{-}$.
@@
    The map $z,\zeta\mapsto z\cdot\zeta$ of ${\cal Z}_{\pm}\times\overline{{\cal C}}_{\mp}$ onto $\bf C$$^{*}$ can be lifted to a map of $\widetilde{\cal Z}_{\pm}\times\overline{{\cal C}}_{\mp}$ onto the covering $\widetilde{\bf C}^{*}$ of the punctured plane $\bf C$$^{*}$. by abuse of language we denote the lifted scalar product by the same symbol.

### v2

@@ 1 Introduction
$W^{(d)}_{\lambda}(z_{1},z_{2})$ $=$ $c_{d}(\lambda)\int_{\gamma(z_{1})}(z_{1}\cdot\zeta)^{\lambda}(z_{2}\cdot\zeta)^{-\lambda-d+1}d\mu_{\gamma}(\zeta)=$ (1.1)
$=$ $c_{d}(\nu)\int_{\gamma(z_{1})}(z_{1}\cdot\zeta)^{-\frac{d-1}{2}+\nu}(z_{2}\cdot\zeta)^{-\frac{d-1}{2}-\nu}d\mu_{\gamma}(\zeta).$ (1.2)
@@
The significance of the new representation (1.1) is evident: it provides the closest possible analogue, within the AdS framework, to the plane-wave Fourier representation of the two-point function for a massive scalar field in (complex) Minkowski space $M_{d}^{(c)}$ with positive energy spectrum [12] which we quote here
@@
and to the plane wave expansion for the de Sitter quantum fields we have constructed long ago [13, 14]. Manifestly covariant momentum space representations of quantum fields provide the physically intuitive ingredients of Feynman diagrammatics [15] and a great deal of the predictive power of quantum field theory depends on them. Furthermore, many crucial mathematical developments such as the PCT and the Spin-Statistics theorems [12] and the Euclidean formulation of quantum field theory [16] are based on the plane-wave Fourier representation of the correlation functions and the positivity of the energy spectrum. The Fourier representation (1.3) and the related representation of the Feynman propagator are of paramount importance. We expect (1.1) to be just as fundamental in the framework of AdS QFT.
The second major topic addressed in this work pertains to the horospherical, or Poincaré, coordinate system, which covers exactly one half of the AdS manifold. We analyze the Fourier transform of the plane waves w.r.t. the Poincaré coordinates and show how this transform diagonalizes the two-point function (1.1). We present here the final result, which exhibits remarkable elegance, while deferring all detailed derivations to Sect. 13 and 14:
@@
Some historical remarks on this integral representation may help to put it in perspective. The Euclidean counterpart of Eq. (1.5), which represents the Schwinger propagator, appears in Bateman’s tables of integral transforms [17, Eq. (12), p. 64]. There, a Legendre function of the second kind [18], with a rather intricate argument, is identified as the Hankel transform of a product of two Bessel functions.
A particular case of Eq. (1.5), corresponding to spacetime dimension $d=2$, is discussed in Watson’s classic treatise [19], together with indications of the proof. A similar integral—Watson notes—had already been discovered by Kirchhoff as early as 1853. Watson also credits other particular cases to [20, 21, 22, 23] and refers to all of them as Lipschitz–Hankel integrals. Formula (1.5) may therefore be called a Lipschitz–Hankel formula [24], or simply a Hankel formula, as we shall do throughout the rest of this paper.
In the AdS literature, the Euclidean Hankel formula for the Schwinger propagator was reported in an appendix [27, Eq. (A.3)], where it appears as an intermediate step toward reconstructing a more conventional representation of the Feynman propagator, always in terms of Bessel functions, previously introduced by [28].
The geometrical interpretation in anti-de Sitter (AdS) space of the Hankel formula (1.5), together with its relation to the maximally analytic Wightman functions of AdS quantum field theory [10]—which also include the Euclidean Schwinger propagator—was established shortly thereafter in [25, 26]. This result was obtained by employing complexified Poincaré coordinates and exploiting the general analyticity properties of Wightman functions that satisfy the AdS spectral condition [10]. From a technical perspective, the Hankel formula was derived in [25] by solving a Schrödinger-type equation in the spacelike coordinate transverse to the Poincaré foliation. A key ingredient in this derivation is Bateman’s formula, whose geometrical interpretation was clarified in that work. This approach also provides insight into the so-called Breitenlohner–Freedman phenomenon [9]. In fact, the operator appearing in the transverse Schrödinger problem is not essentially self-adjoint, and its possible self-adjoint extensions correspond to different admissible quantizations in the regime where two such schemes coexist, thereby offering an alternative interpretation of the Breitenlohner–Freedman phenomenon.
The Lorentzian Hankel formula for the Feynman propagator, which had been somewhat overlooked for a period of time, was eventually revived and brought to general attention in [29, 30]. These studies not only reintroduced the formula but also demonstrated its significant utility in calculations of propagators, correlation functions, and scattering amplitudes in the AdS background. Since then, the formula has become part of the standard AdS toolbox.
Here, the Hankel formula and its untapped potential role in AdS QFT are brought into sharper focus. In particular, its connection to the complex structure of the AdS manifold, the geometry of the complex cone, and the plane-wave solutions of the Klein–Gordon operator is fully elucidated. This perspective highlights the crucial role played by the global analyticity properties of the plane waves, which are inherited from the AdS spectral condition [10], and makes their significance evident in the context of the new derivation.
Moreover, our analysis uncovers an unexpected and intriguing insight. It demonstrates that Euclidean Feynman diagrams constructed in Lobachevsky space can, through a standard Wick rotation, be mapped onto diagrams entirely supported within the Poincaré patch of AdS. Several concrete examples are provided that support the conjecture that any Feynman or Witten diagram, computed by integration solely over a Poincaré patch of the real AdS manifold, continues to respect full AdS covariance, even though the Poincaré patch itself is not invariant under the full isometry group.
@@
Finally in Appendix A we discuss some technical details and in Appendix B we revisit the so called ”split representation” of the Euclidean propagator
@@ 2 AdS and its covering: geometrical setup
The restriction to $AdS_{d}$ of the pseudo-Riemannian metric $g_{\mu\nu}dx^{\mu}dx^{\nu}$ is a Lorentzian metric with signature (+,,,)(+,\ -$\ldots,$-). $G_{0}$ acts transitively on $AdS_{d}$; we chose
@@
to be the distinguished origin of $AdS_{d}$; in the following we will take for simplicity a unit radius $R=1$.
The complex $d$-dimensional anti-de Sitter manifold $AdS_{d}^{(c)}$ can analogously be defined by its embedding in the complex space ${\bf C}^{d+1}_{2}$ endowed with the metric $g_{\mu\nu}$:
$AdS_{d}^{(c)}=\{z=x+iy\in{\bf C}^{d+1}_{2},\ z^{2}=z\cdot z=1\}=\{z=x+iy,\ x^{2}-y^{2}=1,\ x\cdot y=0\}.$ (2.10)
@@
It is possible to single out in $AdS_{d}^{(c)}$ a Euclidean manifold $EAdS_{d}$; we choose the connected real submanifold
$EAdS_{d}=\{z^{\scriptscriptstyle{E}}=(ix^{0},\vec{x},x^{d})\in AdS_{d}^{(c)},\ \ x^{d}\geq 1\}$ (2.11)
equipped with the Riemannian metric induced by the scalar product (2.6). It is the upper sheet of the two-sheeted hyperboloid with equation
${x^{d}}^{2}-{x^{0}}^{2}-{|\vec{x}|}^{2}=1$ (2.12)
i.e. it is a copy of the Lobachevsky space $H_{d}$.
Global coordinates may be introduced on $AdS_{d}$ by the map
$(t,\,\vec{x})\rightarrow\left(\sqrt{1+\vec{x}^{2}}\,\sin t,\ \vec{x},\ \sqrt{1+\vec{x}^{2}}\,\cos t\right)$ (2.13)
@@
$(s,\,\vec{x})\rightarrow(i\sqrt{1+\vec{x}^{2}}\,{\rm sh}\ s,\ \vec{x},\ \sqrt{1+\vec{x}^{2}}\,{\rm ch}\ s).$ (2.14)
@@ 4 The chiral tuboids $\cal Z$$_{+}$ and $\cal Z$$_{-}$
(ii) The image of either of the domains $\cal Z$$_{+}$ or $\cal Z$$_{-}$ under the coordinate map $z\rightarrow z^{d}=e_{d}\cdot z$ is the cut-plane $\Delta={\bf C}\setminus[-1,1]$.
(iii) The image of the domain ${\cal Z}_{-}\times{\cal Z}_{+}$ (or ${\cal Z}_{+}\times{\cal Z}_{-}$) by the scalar product mapping $z_{1},z_{2}\rightarrow z_{1}\cdot z_{2}$ is the cut-plane $\Delta$.
(iv) The map $z_{1},z_{2}\rightarrow z_{1}\cdot z_{2}$ of ${\cal Z}_{-}\times{\cal Z}_{+}$ onto $\Delta$ can be lifted to a map of $\widetilde{\cal Z}_{-}\times\widetilde{\cal Z}_{+}$ onto the covering $\widetilde{\Delta}$ of the cut-plane $\Delta$.
@@ 2. Locality: if $x_{1}$ and $x_{2}$ are space-like separated then the commutator
$4^{\prime\prime}$. Maximal analyticity property: the reduced two-point function $W(z_{1}\cdot z_{2})$ extends to a function analytic in the covering $\widetilde{\Delta}_{1}$ of the cut-plane $\Delta_{1}=\{{\bf C}\setminus[-1,1]\}.$ For theories single-valued on $AdS_{d}$, $W(z_{1}\cdot z_{2})$ is analytic in $\Delta_{1}$.
@@ 6 Klein-Gordon equation, homogeneous functions and plane waves
This remains true if $h$ is a $G_{0}$-invariant distribution such as $\delta((x\cdot x)-1)$. In particular the operators ${\cal M}_{\mu\nu}$ and ${\cal M}^{2}$ commute with the restriction to invariant submanifolds of ${\bf R}^{d+1}_{2}$ and also with dilations.
If $f$ is a smooth function or a distribution homogeneous of degree zero on the open set $\Omega=\{x\in{\bf R}^{d+1}_{2}\ :\ (x,x)>0\}$ then
@@
Eqs. (6.48) and (6.49) then imply that the restriction of $f$ to the anti-de Sitter manifold $AdS_{d}$ satisfies the Klein-Gordon equation
$(\Box+m^{2}_{\lambda})(f(x)|_{AdS_{d}})=0$ (6.50)
@@
Plane waves provide the most important example of homogeneous solutions of the massless equation in the ambient space:
$x,\zeta\rightarrow(x\cdot\zeta)^{\lambda},$ (6.52)
where $\lambda$ is a complex parameter and $\zeta$ is a complex null vector in ${\cal C}_{d}^{(c)}$. The restriction $\phi_{\lambda}(x,\zeta)$ of (6.52) to the anti–de Sitter manifold satisfies the massive Klein–Gordon equation (6.50):
$\phi_{\lambda}(x,\zeta)=(x\cdot\zeta)^{\lambda},\qquad x\in AdS_{d}.$ (6.53)
The invariance of $m_{\lambda}^{2}$ under the involution
$\lambda\to\bar{\lambda}=1-d-\lambda$ (6.54)
implies that, for any given complex null vector $\zeta$, there exists a second solution of Eq. (6.50), obtained by applying this involution to the exponent of the wave:
@@
To establish a connection with notations commonly used in the AdS–CFT literature, we introduce three additional complex parameters: $\nu$, $\Delta_{-}$, and $\Delta_{+}$. These parameters are related to the homogeneity degree $\lambda$ through
$\lambda=-\frac{d-1}{2}+\nu=-\Delta_{-}.$ (6.56)
Under the involution (6.54)
$\bar{\lambda}=1-d-\lambda=-\frac{d-1}{2}-\nu=-\Delta_{+}$ (6.57)
the complex parameter $\nu$ simply changes sign,
$\nu\rightarrow-\nu;$ (6.58)
the involution does not coincide with complex conjugation unless $\nu$ is purely imaginary, which corresponds to a tachyonic wave devoid of physical interpretation in AdS.
We may further clarify the relation between the homogeneity degree $\lambda$ and the scaling dimnsion $\Delta_{-}$ by using the Poincaré coordinates (12.155) and (12.164) to parametrize the plane wave:
$(x(t,u)\cdot\xi(\eta))^{\lambda}=(x(t,u)\cdot\xi(\eta))^{-\frac{d-1}{2}+\nu}=\left(\frac{u}{2}-\frac{\left({{\rm x}}-{\eta}\right)^{2}}{2u}\right)^{-\Delta_{-}}\simeq\frac{u^{\Delta_{-}}}{(\left({{\rm x}}-{\eta}\right)^{2})^{\Delta_{-}}}$ (6.59)
$(x(t,u)\cdot\xi(\eta))^{\bar{\lambda}}=(x(t,u)\cdot\xi(\eta))^{-\frac{d-1}{2}-\nu}=\left(\frac{u}{2}-\frac{\left({{\rm x}}-{\eta}\right)^{2}}{2u}\right)^{-\Delta_{+}}\simeq\frac{u^{\Delta_{+}}}{(\left({{\rm x}}-{\eta}\right)^{2})^{\Delta_{+}}}$ (6.60)
(unless $\rm x$} and $\eta$ are colinear). The squared mass can then be written as
$m_{\lambda}^{2}=\nu^{2}-\frac{(d-1)^{2}}{4}=-\Delta_{+}\Delta_{-}.$ (6.61)
Let us assume that the spacetime dimension parameter $d$ is real. For real $\lambda$ (equivalently, for real $\nu$), the squared mass (6.61) reaches its minimum negative value at $\nu=0$:
$m^{2}_{\min}=-\frac{(d-1)^{2}}{4}.$ (6.62)
This minimum corresponds to the homogeneity degree
$\lambda_{d}=-\frac{d-1}{2}.$ (6.63)
In the following, we will use $\lambda$ and $\nu$ interchangeably.
The main challenge now emerges: although the complex parameter $\lambda$ in the above definition can take any value, the plane waves themselves are only locally well-defined on AdSd. An exception occurs when $\lambda$ is an integer; in this case, the plane waves are globally well-defined. For non-integer $\lambda$, however, constructing single-valued plane waves requires additional effort on the covering space $\widetilde{\text{AdS}}_{d}$. This issue will be discussed in detail in the following section.
@@ 7 Chiral cones
    The image of either $\cal C$$_{+}$ or $\cal C$$_{-}$ under the coordinate map $\zeta\rightarrow\zeta^{d}=e_{d}\cdot\zeta$ is the punctured complex-plane ${\bf C}^{*}={\bf C}\setminus\{0\}$. This fact is intuitively clear as the chiral cones are the limit of the chiral tuboids when the radius $R$ tends to zero.
@@
    Eqs. (7.69) and (7.71) then imply that the image of ${\cal Z}_{-}\times\overline{{\cal C}}_{+}$ by the scalar product mapping $z,\zeta\rightarrow z\cdot\zeta$ is also the punctured complex-plane $\bf C$$^{*}$ and the same is obviously true for ${\cal Z}_{+}\times\overline{{\cal C}}_{-}$. In words: the scalar product $z\cdot\zeta$ cannot vanish when either $z\in{\cal Z}_{-}$ and $\zeta\in\overline{{\cal C}}_{+}$ or when $z\in{\cal Z}_{+}$ and $\zeta\in\overline{{\cal C}}_{-}$.
@@
    The map $z,\zeta\rightarrow z\cdot\zeta$ of ${\cal Z}_{\pm}\times\overline{{\cal C}}_{\mp}$ onto $\bf C$$^{*}$ can be lifted to a map of $\widetilde{\cal Z}_{\pm}\times\overline{{\cal C}}_{\mp}$ onto the covering $\widetilde{\bf C}^{*}$ of the punctured plane $\bf C$$^{*}$. by abuse of language we denote the lifted scalar product by the same symbol.